# Optimal Tests of the Composite Null Hypothesis Arising in Mediation Analysis

### v2

@@ 4 BAYES RISK OPTIMAL TEST
We now consider a Bayes risk optimality criterion. We draw inspiration from and modify the testing procedure of Rosenblum et al., (2014), which gives approximately Bayes risk optimal tests of an entirely different causal hypothesis. They were interested in simultaneously testing for treatment effects both in the overall population and in two subpopulations. Their approach was to discretize the test statistic space into a fine grid of cells, then cast the problem as a constrained optimization problem, where the unknown parameters to optimize over are the rejection probabilities in each cell, resulting in a rejection region that optimizes the Bayes risk they defined specific to their problem. This approach turns out to be easily adapted to our problem of testing the composite null hypothesis $H_{0}$.
We let $M:\mathbb{R}^{2}\times[0,1]\rightarrow\{0,1\}$ denote a generic testing function of $H_{0}$ against $H_{1}$ that characterizes the randomized test consisting of rejecting $H_{0}$ for $H_{1}$ if and only if $M(Z^{x}_{*},Z^{y}_{*},U)=1$, where $U$ is a uniform random variable with support $[0,1]$. As in Rosenblum et al., (2014), it is purely for computational reasons that we consider a randomized test. Doing so yields a linear programming problem rather than an integer programming problem, the latter being far more computationally burdensome than the former, especially for our fairly high-dimensional optimization problem. However, our solution is almost entirely deterministic, with very few cells having non-degenerate rejection probabilities. Let $L:\{0,1\}\times\mathbb{R}^{2}\rightarrow\mathbb{R}$ be a bounded loss function, and $\Lambda$ denote a prior distribution on $(\delta_{x}^{*},\delta_{y}^{*})$. We consider both a 0-1 loss function given by
$L\{M(Z^{x}_{*},Z^{y}_{*},U),\delta_{x}^{*},\delta_{y}^{*}\}=M(Z^{x}_{*},Z^{y}_{*},U)I(\delta_{x}^{*}\delta_{y}^{*}=0)+\{1-M(Z^{x}_{*},Z^{y}_{*},U)\}I(\delta_{x}^{*}\delta_{y}^{*}\neq 0),$
and a (bounded) quadratic loss function given by
$L\{M(Z^{x}_{*},Z^{y}_{*},U),\delta_{x}^{*},\delta_{y}^{*}\}=\{1-M(Z^{x}_{*},Z^{y}_{*},U)\}\{(\delta_{x}^{*}\delta_{y}^{*})^{2}\wedge 4\Phi^{-1}(1-\alpha/2)^{2}\}.$
We consider $\mathcal{N}\{0,2\Phi^{-1}(1-\alpha/2)\bm{I_{2}}\}$ for the prior $\Lambda$. The constrained Bayes optimization problem is the following: For given $\alpha\in(0,1)$, $L$, and $\Lambda$, find the function $M$ minimizing the Bayes risk
@@
$\mathrm{Pr}_{\delta_{x}^{*},\delta_{y}^{*}}\{M(Z^{x}_{*},Z^{y}_{*},U)=1\}\leq\alpha\text{ for all }(\delta_{x}^{*},\delta_{y}^{*})\in\mathbb{R}^{2}\text{ such that }\delta_{x}^{*}\delta_{y}^{*}=0.$ (9)
To formulate an approximation of this problem, we begin by making the test coincide outside of $B:=[-b,b]\times[-b,b]$ with the joint significance test, where we select $b=2\Phi^{-1}(1-\alpha/2)$. The reason for this is that the event $(Z^{x}_{*},Z^{y}_{*})\not\in B$ is only likely under $(\delta_{x}^{*},\delta_{y}^{*})$ far enough away from the origin and very unlikely otherwise, so that the joint significance test will only be negligibly conservative. We then discretize the null hypothesis space into a grid of $18K+$ 1 points and $B$ into a $(2K+1)\times(2K+1)$ grid consisting of squares $R_{k,k^{\prime}}$, where $K$ is a suitably large integer. Within $B$, we define a class of random rejection regions whose realizations are constant within each square. That is, if $(z^{x},z^{y})\in R_{k,k^{\prime}}$, $M(z^{x},z^{y},u)=M(z^{x^{\prime}},z^{y^{\prime}},u)$ for all $(z^{x^{\prime}},z^{y^{\prime}})\in R_{k,k^{\prime}}$ and all $u\in[0,1]$. The rejection probabilities corresponding to each square are the unknown arguments defining the rejection region that are to be optimized over. A discretized version of the constrained optimization problem is then defined corresponding to these grids. Collectively, these form a linear optimization problem that approximates the Bayes risk optimization problem. Full details of the discretization and corresponding approximation to the optimization problem are available in the Supplementary Materials. The linear optimization problem is high-dimensional, with the dimension of the unknowns being $4K^{2}$, and the number of constraints being $18K^{2}+8K+$ 1. However, this is also a highly sparse optimization problem, which allows for fairly efficient computation despite the dimensionality. The type 1 error constraints consist of only $18K+$ 1 inequalities, and the matrix characterizing the remaining probability constraints contains only one nonzero element per row.
@@
For both tests, the rejection probabilities are nearly all zero or one. If one prefers a nonrandom test, one can obtain a slightly more conservative test by setting all nondegenerate probabilities to zero. Although we did not enforce symmetry in this particular implementation of the optimization problem, the plots are nonetheless almost completely symmetrical. Each test almost contains the entire rejection region of the joint significance test. Both rejection regions take interesting, somewhat unexpected shapes, with four roughly-symmetric disconnected regions inside of the acceptance region in the case of the 0-1 loss, and a single region inside the acceptance region centered around (0,0) in the case of the quadratic loss. As with the minimax optimal test, it seems that it is important to have part of the rejection region along the diagonals at least somewhat close to the origin, so that smaller alternatives in both $\delta_{x}^{*}$ and $\delta_{y}^{*}$ have larger rejection probabilities to offset the conservativeness induced by lying in the intersection of two acceptance region bands. As with the minimax optimal test, one might take issue with rejecting test statistics that are arbitrarily close to the null hypothesis space. We discuss this further in Section 6.
While solving the linear programming problem to produce this test takes a nontrivial amount of time to run, this need only be done once, after which the object is stored and can be quickly loaded on demand. The approximate Bayes risk optimal tests can be solved for any size $\alpha\in(0,1)$. However, the linear optimization problem must be solved separately for each value of $\alpha$, and so cannot be obtained immediately.
As with the minimax optimal test, the approximate optimality of the tests defined in this section is in terms of the limiting distribution $(Z_{*}^{x},Z_{*}^{y})$. Assuming uniform convergence of $(\hat{\delta}_{x},\hat{\delta}_{y})$ as in (1), the (sample) approximations to the (computational) approximate Bayes risk optimality objective functions and type 1 error constraints can be made arbitrarily small uniformly over the parameter space given a sufficiently large sample size. As with the minimax optimal test, the asymptotic approximation can be improved by replacing all instances of the normal distribution with the $t$ distribution with $(n-1)$ degrees of freedom.
@@ 5 LARGE-SCALE HYPOTHESIS TESTING
As mentioned above, applying the Bonferroni correction to the proposed methods in order to control for familywise error rate is straightforward even without the use of $p$-values. For $J$ hypotheses, one can simply run any of the proposed tests with level $\alpha/J$ instead of comparing the $p$-value to a threshold of $\alpha/J$. Adapting the Benjamini–Hochberg procedure to control false discovery rate requires a bit more explanation if $p$-values are not available (though the $p$-value proposed in the Supplementary Materials can be used, which will be conservative). A modified Benjamini–Hochberg procedure works as follows: (i) find the largest value $j$ such that at least $j$ of the $J$ tests reject at $\alpha$ level $\alpha j/J$, (ii) reject all hypotheses that the test rejects at level $\alpha j/J$. It can be shown that this procedure controls the false discovery rate at $\alpha$ even for tests that are not monotone in $\alpha$. A (potentially) conservative version of this procedure can avoid an exhaustive search for the largest $j$ in step (i). This procedure starts with $j=1$ and increases $j$ until fewer than $\alpha j/J$ hypotheses are rejected for $K$ consecutive iterations for some pre-specified $K$, after which the largest $j$ in this sequence at which at least $\alpha j/J$ hypotheses are rejected is chosen for part (i).
One important and often overlooked point when testing for the presence of multiple mediated effects is that one needs to be willing to assume that the true mediators do not affect one another, which would result in some mediators being exposure-induced confounders, yielding the NIE non-identifiable (Avin et al.,, 2005). This may not be realistic in many scenarios; however, this problem is beyond the scope of this article.
@@ 6 INTERPRETATION AND MODIFICATIONS
As we have pointed out, the rejection regions for the minimax and Bayes risk optimal tests have some peculiar features that some may regard as undesirable. Tests with such properties are known to arise in multi-parameter hypothesis testing and have received criticism in Perlman and Wu, (1999) as being flawed, though the philosophical debate about what criteria should be used to compare among statistical tests is perhaps not entirely settled (Berger,, 1999; McDermott and Wang,, 1999). Rather than choosing a side on this debate, take a pluralistic stance and provide a menu of options, among which we hope anyone across this ideological spectrum can find a test that suits their preferences. Furthermore, we discuss differences in interpretation of different tests.
@@
Not all of the criticisms of Perlman and Wu, (1999) apply to our tests, but at least two do. First, the minimax optimal test can reject when the point estimate is arbitrarily close to the null. How then is one to interpret a test that rejects when both test statistics are very small? One perspective is that a scientific investigator need not rely entirely on the result of a hypothesis test to determine whether an association or an effect is meaningful or “practically significant”, so long as the hypothesis test has correct type 1 error. Indeed, one could always pre-specify a minimally meaningful effect size, and never reject if the point estimate does not exceed it.
Second, the minimax optimal test is not monotonic in $\alpha$, nor is it guaranteed that the Bayes risk optimal test is. In the Supplementary Materials, we introduce a $p$-value corresponding to the minimax optimal test, which does not correspond one-to-one with the minimax optimal test. In fact, the $p$-value can itself be used to define a new test that rejects at level $\alpha$ whenever the $p$-value is less than $\alpha$. A constrained variant of the Bayes risk optimal test can be devised for a given sequence of $\alpha$ values. This can be done by setting the rejection value to be no less (greater) than the largest (smallest) rejection probability among all tests with larger (smaller) $\alpha$ values preceding it in the sequence. Furthermore, the same realization of $U$ ought to be used across tests to preserve monotonicity.
@@
To circumvent the first issue, we propose modifications of the minimax optimal and Bayes risk optimal tests. For the minimax optimal test and its corresponding $p$-value-based test, one can simply truncate the rejection region according to any user-specified distance, say $d$, from the null hypothesis space. For a test with rejection region $R$, the rejection region of the corresponding truncated version will simply be $R\setminus\{(Z_{n}^{x},Z_{n}^{y})^{\top}:\lvert Z_{n}^{x}\rvert\wedge\lvert Z_{n}^{y}\rvert\leq d\}$. The rejection region of the truncated minimax optimal test with $d=0.1$ is shown in panel (a) of Figure 3. Of course, any cutoff can be chosen; we simply illustrate with 0.1 for the sake of comparison, because this is the distance of the rejection region of van Garderen and van Giersbergen, 2022a from the null hypothesis space.
@@
Such truncated tests will clearly be dominated by their corresponding non-truncated versions in terms of power. A more powerful approach would be to instead optimize a constrained version of the minimax optimality criterion, where the constraint is that the test cannot reject for test statistics such that $\lvert Z_{n}^{x}\rvert\wedge\lvert Z_{n}^{y}\rvert\leq d$. This is a challenging problem that we do not address here, but that could be a fruitful direction for future research. One rebuttal to the claim of Perlman and Wu, (1999) that the rejection region ought to be bounded away from the null hypothesis space is that this distance is not specified, nor is there a clearly-defined criterion specifying what this distance should be (Berger,, 1999). However, if one believes that this distance should be determined according to the likelihood ratio test, then one would set $d$ to be $\Phi^{-1}(1-\alpha/2)$, in which case the joint significance test, i.e., the likelihood ratio test (Liu et al.,, 2022; van Garderen and van Giersbergen, 2022a, ), would be recovered in the single hypothesis test case, and our test offers nothing novel. However, if one is conducting multiple hypothesis testing and wishes to control for familywise error rate (for instance), and insists on a $d$ of $\Phi^{-1}(1-\alpha/2)$, then this does indeed generate a novel test with greatly-improved power over the Bonferroni-corrected joint significance test. For example, for a Bonferroni correction for 100 tests at level $\alpha=0.05$, one can use the truncated minimax optimal test with $\alpha$ level $0.05/100$ and with $d=\Phi^{-1}(1-\alpha/2)$, which yields the rejection region in panel (b) of Figure 3. In other words, one can apply the unmodified, but multiple testing-adjusted minimax optimal test, and then use the non-multiple testing-adjusted joint significance test as an additional test to screen out effects deemed too small. In this case, the Bonferroni-corrected minimax optimal test will have 100 times the worst-case power per test of that of the Bonferroni-corrected joint significance test, while still preserving familywise error rate and respecting the specified bound.
@@
Another important challenge has to do with the traditional interpretation of a hypothesis test rejection decision: if the null hypothesis were true, then the probability of observing an event as or more extreme than the observed event is at most $\alpha$. In multi-parameter hypothesis testing settings, this interpretation can be ambiguous, as there is not a total ordering in the parameter space. Nevertheless, in the product of coefficients case considered here, such an interpretation can be applied to a partial ordering, where one pair of test statistics is “as or more extreme” than another if both elements in the first pair are at least as large in magnitude as those in the second. van Garderen and van Giersbergen, 2022b refer to a test that rejects for all pairs of test statistics that are more extreme (in this sense) than another pair for which it also rejects as “information coherent”. Clearly none of the tests we have proposed nor the test of van Garderen and van Giersbergen, 2022a satisfy this property, hence these tests lack this traditional interpretation. Instead, they admit a weaker interpretation: if the null hypothesis were true, then the observed event would have been unlikely to occur in the sense that its rejection probability would be at most $\alpha$. On the other hand, if the joint significance test and one of our proposed tests both reject, then the stronger interpretation can be applied. In a case where they disagree, then if one is applying one of our proposed tests, then one must be prepared to accept the possibility that one pair of test statistics that are smaller in magnitude may reject while another pair does not; this is entirely consistent with this weaker interpretation.
van Garderen and van Giersbergen, 2022b showed that the joint significance test is the most powerful test in the class of information coherent tests. Thus, there is a fundamental trade-off between gaining power over the joint significance test and preserving the former stronger interpretation over the latter weaker interpretation. If one deems the underpoweredness of the joint significance test to be an important problem, then one has no choice but to sacrifice a degree of interpretability. In proposing more powerful tests, we wish to present options to gain power at the cost of some interpretability for those who would deem it a worthwhile trade-off. One may reasonably feel uncomfortable making such a trade-off when scientific conclusions or policy decisions are to be based on the result. However, one may feel more comfortable with this trade-off in multiple hypothesis testing scenarios where the focus is on making multiple discoveries while controlling the number of errors. Thus, our proposed methodology is perhaps best suited for this latter setting.

### v3

@@ 4 BAYES RISK OPTIMAL TEST
We now consider a Bayes risk optimality criterion. We draw inspiration from and modify the testing procedure of Rosenblum et al., (2014), which gives approximately Bayes risk optimal tests for the distinct purpose of simultaneously testing for treatment effects both in the overall population and in two subpopulations. Their approach was to discretize the test statistic space into a fine grid of cells, then cast the problem as a constrained optimization problem, where the unknown parameters are the rejection probabilities in each cell, resulting in a rejection region that optimizes their particular Bayes risk.
We let $M:\mathbb{R}^{2}\times[0,1]\rightarrow\{0,1\}$ denote a generic testing function of $H_{0}$ against $H_{1}$ that characterizes the randomized test consisting of rejecting $H_{0}$ for $H_{1}$ if and only if $M(Z^{x}_{*},Z^{y}_{*},U)=1$, where $U$ is a uniform random variable on $[0,1]$. As in Rosenblum et al., (2014), it is purely for computational reasons that we consider a randomized test, as doing so yields a linear programming problem rather than an integer programming problem. However, our solution is almost entirely deterministic, with very few cells having non-degenerate rejection probabilities. Let $L:\{0,1\}\times\mathbb{R}^{2}\rightarrow\mathbb{R}$ be a bounded loss function, and $\Lambda$ denote a prior distribution on $\bm{\delta}^{*}$. We consider a 0-1 loss function given by
$L\{M(Z^{x}_{*},Z^{y}_{*},U),\bm{\delta}^{*}\}=M(Z^{x}_{*},Z^{y}_{*},U)I(\delta_{x}^{*}\delta_{y}^{*}=0)+\{1-M(Z^{x}_{*},Z^{y}_{*},U)\}I(\delta_{x}^{*}\delta_{y}^{*}\neq 0),$
and a bounded quadratic loss function given by
$L\{M(Z^{x}_{*},Z^{y}_{*},U),\bm{\delta}^{*}\}=\{1-M(Z^{x}_{*},Z^{y}_{*},U)\}\{(\delta_{x}^{*}\delta_{y}^{*})^{2}\wedge 4\Phi^{-1}(1-\alpha/2)^{2}\}.$
We consider $\mathcal{N}\{0,2\Phi^{-1}(1-\alpha/2)\bm{I_{2}}\}$ for the prior $\Lambda$. The constrained Bayes optimization problem is the following: For given $\alpha\in(0,1)$, $L$, and $\Lambda$, find the function $M$ minimizing
@@
$\mathrm{Pr}_{\bm{\delta}^{*}}\{M(Z^{x}_{*},Z^{y}_{*},U)=1\}\leq\alpha\text{ for all }\bm{\delta}^{*}\in\mathbb{R}^{2}\text{ such that }\delta_{x}^{*}\delta_{y}^{*}=0.$ (9)
To formulate an approximation of this problem, we begin by making the test coincide outside of $B:=[-b,b]\times[-b,b]$ with the joint significance test, where we select $b=2\Phi^{-1}(1-\alpha/2)$. This is because the event $(Z^{x}_{*},Z^{y}_{*})\not\in B$ is only likely under $\bm{\delta}^{*}$ far enough away from the origin and very unlikely otherwise, so that the joint significance test will only be negligibly conservative. We then discretize the null hypothesis space into a grid of $18K+$ 1 points and $B$ into a $(2K+1)\times(2K+1)$ grid consisting of squares $R_{k,k^{\prime}}$, where $K$ is a suitably large integer. Within $B$, we define a class of random rejection regions whose realizations are constant within each square. That is, if $(z^{x},z^{y})\in R_{k,k^{\prime}}$, then $M(z^{x},z^{y},u)=M(z^{x^{\prime}},z^{y^{\prime}},u)$ for all $(z^{x^{\prime}},z^{y^{\prime}})\in R_{k,k^{\prime}}$ and all $u\in[0,1]$. The rejection probabilities corresponding to each square are the unknown arguments defining the rejection region that are to be optimized over. A discretized version of the constrained optimization problem is then defined corresponding to these grids. Collectively, these form a linear optimization problem that approximates the Bayes risk optimization problem. Full details are available in the Supplementary Materials.
@@
For both tests, the rejection probabilities are nearly all zero or one. If one prefers a nonrandom test, one can obtain a slightly more conservative test by setting all nondegenerate probabilities to zero. Both rejection regions almost contain the entire rejection region of the joint significance test. Moreover, both rejection regions take interesting, somewhat unexpected shapes, with four roughly-symmetric disconnected regions inside of the acceptance region in the case of the 0-1 loss, and a single region inside the acceptance region centered around (0,0) in the case of the quadratic loss. As with the minimax optimal test, it seems that it is important to have part of the rejection region along the diagonals at least somewhat close to the origin, so that smaller alternatives in both $\delta_{x}^{*}$ and $\delta_{y}^{*}$ have larger rejection probabilities to offset the conservativeness induced by lying in the intersection of two acceptance region bands. As with the minimax optimal test, one might reasonably take issue with rejecting test statistics that are arbitrarily close to the null hypothesis space. We discuss this further in Section 6.
While solving the linear programming problem to produce these tests takes a nontrivial amount of time to run, this need only be done once, after which the object is stored and can be quickly loaded on demand. The approximate Bayes risk optimal tests can be solved for any size $\alpha\in(0,1)$. However, the linear optimization problem must be solved separately for each value of $\alpha$, and so cannot be obtained immediately.
As with the minimax optimal test, the approximate optimality of the tests defined in this section is in terms of the limit distribution, i.e., of $(Z_{*}^{x},Z_{*}^{y})$. Assuming uniform convergence of $(\hat{\delta}_{x},\hat{\delta}_{y})$ as in (1), the (sample) approximations to the (computational) approximate Bayes risk optimality objective functions and type 1 error constraints can be made arbitrarily small uniformly over the parameter space given a sufficiently large sample size. As with the minimax optimal test, the asymptotic approximation can be improved by replacing all instances of the normal distribution with the $t$ distribution with $(n-1)$ degrees of freedom.
@@ 5 LARGE-SCALE HYPOTHESIS TESTING
As mentioned above, applying the Bonferroni correction to the proposed methods in order to control for familywise error rate is straightforward even without the use of $p$-values. For $J$ hypotheses, one can simply run any of the proposed tests with level $\alpha/J$ instead of comparing the $p$-value to a threshold of $\alpha/J$. Adapting the Benjamini–Hochberg procedure to control false discovery rate requires a bit more explanation if $p$-values are not available (though the $p$-value proposed in the Supplementary Materials can be used, which will be conservative). A modified Benjamini–Hochberg procedure works as follows: (i) find the largest value $j$ such that at least $j$ of the $J$ tests reject at level $\alpha j/J$, (ii) reject all hypotheses that the test rejects at level $\alpha j/J$. It can be shown that this procedure controls the false discovery rate at $\alpha$ even for tests that are not monotone in $\alpha$. A (potentially) conservative version of this procedure can avoid an exhaustive search for the largest $j$ in step (i). This procedure starts with $j=1$ and increases $j$ until fewer than $\alpha j/J$ hypotheses are rejected for $\kappa$ consecutive iterations for some pre-specified $\kappa$, after which the largest $j$ in this sequence at which at least $\alpha j/J$ hypotheses are rejected is chosen for part (i).
When testing for the presence of multiple mediated effects, one needs to assume that the true mediators do not affect one another, which would result in some mediators being exposure-induced confounders, yielding the NIE non-identifiable. This may not be realistic in many scenarios; however, this problem is beyond the scope of this article.
@@ 6 INTERPRETATION AND MODIFICATIONS
As we have pointed out, the rejection regions for the minimax and Bayes risk optimal tests have some peculiar features that some may regard as undesirable. Tests with such properties are known to arise in multi-parameter hypothesis testing and have received criticism in Perlman and Wu, (1999) as being flawed, though the philosophical debate about what criteria should be used to compare among statistical tests is perhaps not entirely settled (Berger,, 1999). We take a pluralistic stance and provide a menu of options, among which we hope anyone across this ideological spectrum can find a test that suits their preferences. Furthermore, we discuss differences in interpretation of different tests.
@@
Not all of the criticisms of Perlman and Wu, (1999) apply to our tests, but at least two do. First, the minimax optimal test can reject when the point estimate is arbitrarily close to the null. How then is one to interpret a test that rejects when both test statistics are small? One perspective is that an investigator need not rely entirely on the result of a hypothesis test to determine whether an association or an effect is “practically significant”, so long as the hypothesis test has correct type 1 error. Indeed, one could always pre-specify either a minimally meaningful effect size or proportion mediated (the natural indirect effect divided by the total effect), and never reject if the point estimate does not exceed it.
Second, the minimax optimal test is not monotonic in $\alpha$, nor is the Bayes risk optimal test guaranteed to be. Thus, in the large-scale hypothesis testing case, it is possible for the Bonferroni- or BH-adjusted minimax optimal test to reject based on a particular test statistic when there are, say, $J$ hypotheses to test, but not when there are $(J-1)$. While this is true for a fixed test statistic, the expected number of rejected nulls will always be lower with a smaller number of hypotheses, and FWER or FDR are preserved for any number of hypotheses. In the context of large-scale hypothesis testing, one might care less about the interpretation of the rejection of a particular hypothesis, and more about minimizing false positives, as rejected nulls may be viewed as being flagged for further scrutiny in subsequent analyses. However, these are not the only options. In the Supplementary Materials, we introduce a $p$-value corresponding to the minimax optimal test, which can itself be used to define a test that rejects at level $\alpha$ whenever the $p$-value is less than $\alpha$. Such a test will be monotone in $\alpha$ by construction. A constrained variant of the Bayes risk optimal test can be devised for a given decreasing sequence of $\alpha$’s by setting the rejection value to be no less than the largest rejection probability among all tests with larger $\alpha$ values preceding it in the sequence. The same realization of the randomization variable $U$ ought to be used across tests to preserve monotonicity. Therefore, if one is concerned about coherence, then one might opt to use this $p$-value based test at the cost of power, but which is still more powerful than traditional tests while preserving FWER or FDR, or alternatively this constrained version of the Bayes risk optimal test.
@@
To circumvent the first issue, we propose modifications of the minimax optimal and Bayes risk optimal tests. For the minimax optimal test, one can simply truncate the rejection region according to any user-specified distance, say $d$, from the null hypothesis space. If one wishes to preserve monotonicity in $\alpha$, then one can do the same for the $p$-value-based test we have just discussed. For a test with rejection region $R$, the rejection region of the corresponding truncated version will simply be $R\setminus\{(Z_{n}^{x},Z_{n}^{y})^{\top}:\lvert Z_{n}^{x}\rvert\wedge\lvert Z_{n}^{y}\rvert\leq d\}$. The rejection region of the truncated minimax optimal test with $d=0.1$ is shown in panel (a) of Figure 3, where we chose 0.1 for the sake of comparison with van Garderen and van Giersbergen, (2024).
@@
Such truncated tests will clearly be dominated by their corresponding non-truncated versions in terms of power. A more powerful approach would be to instead optimize a constrained version of the minimax optimality criterion, where the constraint is that the test cannot reject for test statistics such that $\lvert Z_{n}^{x}\rvert\wedge\lvert Z_{n}^{y}\rvert\leq d$. This is a challenging problem that we do not address here. One rebuttal to the claim of Perlman and Wu, (1999) that the rejection region ought to be bounded away from the null hypothesis space is that this distance is not specified, nor is there a clearly-defined criterion specifying what this distance should be (Berger,, 1999). If one believes that this distance should be determined according to the likelihood ratio test, then one would set $d$ to be $\Phi^{-1}(1-\alpha/2)$, in which case the joint significance test, i.e., the likelihood ratio test (Liu et al.,, 2022; van Garderen and van Giersbergen,, 2024), would be recovered in the single hypothesis test case. However, if one is conducting large-scale hypothesis testing and wishes to control for familywise error rate (for instance), and insists on a $d$ of $\Phi^{-1}(1-\alpha/2)$, then this does indeed generate a novel test with greatly-improved power over the Bonferroni-corrected joint significance test. For example, for a Bonferroni correction for 100 tests at level $\alpha=0.05$, one can use the truncated minimax optimal test with $\alpha$ level $0.05/100$ and with $d=\Phi^{-1}(1-\alpha/2)$, which yields the rejection region in panel (b) of Figure 3. That is, one can apply the multiple testing-adjusted minimax optimal test with level $0.0005$, and then use the joint significance test with $\alpha$ level $0.05$ to screen out effects deemed too small. The resulting test will have 100 times the worst-case power per test of that of the Bonferroni-corrected joint significance test, while still preserving familywise error rate and respecting the specified bound.
@@
Another important challenge has to do with the traditional interpretation of a hypothesis test rejection decision: if the null hypothesis were true, then the probability of observing an event as or more extreme than the observed event is at most $\alpha$. In multi-parameter hypothesis testing settings, this interpretation can be ambiguous, as there is not a total ordering in the parameter space. Nevertheless, in the product of coefficients case, such an interpretation can be applied to a partial ordering, where one pair of test statistics is “as or more extreme” than another if both elements in the first pair are at least as large in magnitude as those in the second. van Garderen and van Giersbergen, (2022) refer to a test that rejects for all pairs of test statistics that are more extreme than another pair for which it also rejects as “information coherent”. Clearly none of the tests we have proposed nor the test of van Garderen and van Giersbergen, (2024) satisfy this property, hence these tests lack this traditional interpretation. Instead, they admit a weaker interpretation: if the null hypothesis were true, then the observed event would have been unlikely to occur in the sense that its rejection probability would be at most $\alpha$. Alternatively, if the joint significance test and one of our proposed tests both reject, then the stronger interpretation can be applied. In a case where they disagree and one is applying one of our proposed tests, then one must be prepared to accept the possibility that one pair of test statistics that are smaller in magnitude may reject while another pair does not.
van Garderen and van Giersbergen, (2022) showed that the joint significance test is the most powerful test in the class of information coherent tests. Thus, there is a fundamental trade-off between gaining power over the joint significance test and preserving the former stronger interpretation over the latter weaker interpretation. If one deems the underpoweredness of the joint significance test to be an important problem, then one has no choice but to sacrifice a degree of interpretability. In proposing more powerful tests, we wish to present options to gain power at the cost of some interpretability for those who would deem it a worthwhile trade-off. One may reasonably feel uncomfortable making such a trade-off when scientific conclusions or policy decisions are to be based on the result. However, one may feel more comfortable with this trade-off in large-scale hypothesis testing scenarios where the focus is on making multiple discoveries while controlling the number of errors. Thus, our proposed methodology is perhaps best suited for this latter setting.